### v4

# Direct finiteness of representable regular *-rings

Abstract.
We show that a von Neumann regular ring with involution is directly finite provided that it admits a representation as a *-ring of endomorphisms of a vector space endowed with a non-degenerate orthosymmetric sesquilinear form.
Key words and phrases:
Regular ring with involution, representation, direct finiteness
1991 Mathematics Subject Classification:
16E50, 16W10
0. Erratum
There is no proof of Thm 3 since there is no proof of $et=er$ in Lemma 4. Actually, there is a counterexample given is the new paper: Direct finiteness of representable regular *-rings: A counterexample.
1. Introduction
A *-ring, that is ring with involution, is called finite if $rr^{*}=1$ implies $r^{*}r=1$. This is a basic notion in the classification of von Neumann algebras; in particular, as shown by Murray and von Neumann, a finite von Neumann algebra admits a finite *-ring of quotients. This ring is also *-regular and directly finite, that is $rs=1$ implies $sr=1$. As Ara and Menal [1] have shown, any *-regular ring is at least finite, while direct finiteness remains an open question, as stated by Handelman [5, Problem 48]. The present note gives a positive answer for certain [von Neumann] regular rings with involution.
For $\ast$-rings, there is a natural and well established concept of [faithful] representation in a vector space $V_{F}$ endowed with a non-degenerate orthosymmetric sesquilinear form: an embedding into the *-ring $\operatorname{{\sf End}^{\ast}}(V_{F})$ of those endomorphisms of $V_{F}$ which admit an adjoint. Famous examples are due to Gel’fand-Naimark-Segal ($C^{\ast}$-algebras in Hilbert space) and Kaplansky (primitive $\ast$-rings with a minimal right ideal). For $\ast$-regular rings of classical quotients of finite Rickart $C^{\ast}$-algebras existence of representations has been established in [8], jointly with M. Semenova. N. Niemann [11, 6] has shown that a subdirectly irredcucible *-regular ring is representable if and only its ortholattice lattice of principal right ideals is representable within the ortholattice of closed subspaces of some $V_{F}$.
According to joint work with Susan M. Roddy [7], representability of modular ortholattices is equivalent to membership in a variety generated by finite height members. Using ideas from Tyukavkin [12], the analogue for *-regular rings was obtained by F. Micol [10]. Here, we rely on the presentation given in [9]: A regular *-ring can be represented within $V_{F}$, respectively some ultrapower thereof, if and only if it can be obtained via formation of ultraproducts, regular *-subrings, and homomorphic images from the class of the $\operatorname{{\sf End}^{\ast}}(U_{F})$, $U_{F}$ ranging over finite dimensional non-degenerate subspaces of $V_{F}$. It will be shown that direct finiteness is inherited under the particular construction of [9] which proves the reduction to finite dimensions.
Thanks are due to Ken Goodearl for the hint to reference [4].
2. Preliminaries
When mentioning rings, we always mean associative rings $R$ with unit $1_{R}$, considered as a constant. In any ring, if $a$ has a left inverse $x$, that is $xa=1$, then $a$ is left cancellable, that is $ay=az$ implies $y=z$. $R$ is directly finite if, for all $r,s\in R$, $sr=1$ implies $rs=1$. In such ring, $a$ has a left inverse if and only if $a$ is a unit (and $x=a^{-1}$). The endomorphism ring $\operatorname{{\sf End}}(V_{F})$ of a vector space is directly finite if and only if $\dim V_{F}<\omega$. A $\ast$-ring is a ring endowed with an involution $r\mapsto r^{*}$; an element $e$ of such a ring is a projection, if $e=e^{2}=e^{\ast}$.
A ring $R$ is [von Neumann] regular if for any $a\in R$, there is an element $x\in R$ such that $axa=a$; such an element is called a quasi-inverse of $a$. If, for all $a$, $x$ can be chosen a unit, then $R$ is unit regular. Examples of such are the $\operatorname{{\sf End}}(V_{F})$, $\dim V_{F}<\omega$. A detailed discussion of direct finiteness in regular rings is given in Goodearl [5]. A regular *-ring is *-regular if $xx^{*}=0$ only for $x=0$.
In a regular ring, any left cancellable $a$ has a left inverse; indeed $axa=a$ implies $xa=1$. If $xa=1$ and $aua=a$ with a unit $u$ then $a=u^{-1}$ and $x=a^{-1}$. It follows
Fact 1.
In a directly finite regular ring every left cancellable element is a unit – similarly on the right. Every unit regular ring is directly finite.
[…]
Consider a linear subspace $U$ of $V_{F}$, $1<\dim U_{F}<\omega$. With the induced sesquilinear form, $U_{F}$ is pre-hermitean if and only if $V=U\oplus U^{\perp}$; in particular, there is a projection $\pi_{U}\in\operatorname{{\sf End}^{\ast}}(V_{F})$ such that $U=\operatorname{{\sf im}}\pi_{U}$ and such that the inclusion map $\varepsilon_{U}:U\to V$ is the adjoint of $\pi_{U}$ (here, considerd as a map $V\to U$). We write in this case $U\in\mathbb{O}(V_{F})$ and say that $U$ is a finite-dimensional orthogonal summand. A crucial fact is that $V_{F}$ is the directed union of the $U_{F}$, $U\in\mathbb{O}(V_{F})$. Let $C(F)$ denote the center of $F$ and, for $U\in\mathbb{O}(V_{F})$,
$\begin{array}[]{lcl}B_{U}&=&\{\varepsilon_{U}\varphi\pi_{U}+\lambda\operatorname{{\sf id}}_{V}\mid\varphi\in\operatorname{{\sf End}^{\ast}}(U_{F}),\ \lambda\in C(F)\}\\ &=&\{\psi\in\operatorname{{\sf End}}(V_{F})\mid\psi(U)\subseteq U\;\&\;\exists\lambda\in C(F)\,\psi|U^{\perp}=\lambda\operatorname{{\sf id}}_{U^{\perp}}\}\end{array}$
($\varphi$ and $\psi$ are related via $\varphi(v)=\psi(v)+\lambda v$). Thus, $B_{U}$ is a $\ast$-subring of $\operatorname{{\sf End}^{\ast}}(V_{F})$ and embeds into $\operatorname{{\sf End}^{\ast}}(W_{F})$ for any $W\in\mathbb{O}(V_{F})$, $U\subset W\neq U$. In particular, $B_{U}$ is directly finite. Moreover, $B_{U}$ is unit regular; indeed, $\chi\in B_{U}$ is a unit quasi-inverse of $\psi\in B_{U}$ if $\chi|U$ is one of $\psi|U$ and $\chi|U^{\perp}=(\psi|U^{\perp})^{-1}$ (considering these as endomorphisms of $U$ and $U^{\perp}$, respectively). We put
$\begin{array}[]{lcl}J(V_{F})&=&\{\varphi\in\operatorname{{\sf End}^{\ast}}(V_{F})\mid\dim\operatorname{{\sf im}}\varphi<\omega\}\\ \hat{J}(V_{F})&=&\{\varphi+\lambda\operatorname{{\sf id}}_{V}\mid\varphi\in J(V_{F}),\,\lambda\in C(F)\}\end{array}$
According to [9, Proposition 4.4] $\hat{J}(V_{F})$ is a $\ast$-subring of $\operatorname{{\sf End}^{\ast}}(V_{F})$ and $J(V_{F})$ is an ideal of $\operatorname{{\sf End}^{\ast}}(V_{F})$ closed under the involution. Also, the following holds.
    For any finite $\Phi\subseteq J(V_{F})$ there is $U\in\mathbb{O}(V_{F})$ such that $\varphi=\pi_{U}\varphi=\varphi\pi_{U}$ for all $\varphi\in\Phi$.
Thus, $\hat{J}(V_{F})$ is the directed union of the $B_{U},\,U\in\mathbb{O}(V_{F})$, whence unit-regular.
Lemma 2.
Every regular $\ast$-subring $R$ of $\operatorname{{\sf End}^{\ast}}(V_{F})$ extends to a regular $\ast$-subring $\hat{R}$ of $\operatorname{{\sf End}^{\ast}}(V_{F})$ containing $\hat{J}(V_{F})$ and such that $J(V_{F})$ is an ideal of $\hat{R}$.
Proof.
$\{\lambda\varphi\mid\varphi\in R,\,\lambda\in C(F)\}$ is a regular $\ast$-subring $R^{\prime}$ of $\operatorname{{\sf End}^{\ast}}(V_{F})$ and [9, Proposition 4.5] applies to $R^{\prime}$. ∎
Recall that a [faithful] representation of a $\ast$-ring $R$ within a pre-hermitian space $V_{F}$ is an embedding $\varepsilon\colon R\to\operatorname{{\sf End}^{\ast}}(V_{F})$. It is convenient to consider representations as unitary $R$-$F$-bimodules ${}_{R}V_{F}$ (where the action of $R$ is given as $rv=\varepsilon(r)(v)$) with sesquilinear form on $V_{F}$; that is, a $3$-sorted structure with sorts $V$, $R$, and $F$. Considering a *-subring $A$ of $R$ we may add a fourth sort, $A$, and the embedding map. to obtain $(_{R}V_{F};A)$. Any elementary extension $(_{\tilde{R}}\tilde{V}_{\tilde{F}};\tilde{A})$ is again such a structure, that is, a representation of $\tilde{R}$ and a *-ring $\tilde{A}$ which may be considered as *-subring of $\tilde{R}$. It is a modestly saturated extension if, for each set $\Sigma(\bar{x})$ of first order formulas in finitely many [sorted] variables and with parameters from $(_{R}V_{F};A)$, one has $(_{\tilde{R}}\tilde{V}_{\tilde{F}};\tilde{A})\models\exists\bar{x}.\Sigma(\bar{x})$, provided that $\Sigma(\bar{x})$ is finitely realized in $(_{R}V_{F};A)$, that is $(_{R}V_{F};A)\models\exists\bar{x}.\Psi(\bar{x})$ for every finite subset $\Psi(\bar{x})$ of $\Sigma(\bar{x})$. Such extension always exists, cf. [2, Corollary 4.3.1.4].
3. Main result
Theorem 3.
Every representable regular $\ast$-ring is directly finite.
Proof.
We recall the relevant steps of the proof of [9, Theorem 10.1]. Given a representation ${}_{R}V_{F}$ of the regular *-ring $R$, we may assume that $\dim V_{F}\geqslant\omega$. In view of Proposition 2, we also may assume that $R$ is a *-subring of $\operatorname{{\sf End}^{\ast}}(V_{F})$ containing $A=\hat{J}(V_{F})$ and having ideal $J(V_{F})$. Choose $(_{\tilde{R}}\tilde{V}_{\tilde{F}};\tilde{A})$ a modestly saturated elementary extension of $(_{R}V_{F};A)$.
Let $J_{0}$ denote the set of projections in $J(V_{F})$. For $a\in\tilde{A}$ and $r\in R$, we put $a\sim r$ if $ae=re$ and $a^{\ast}e=r^{\ast}e$ for all $e\in J_{0}$. According to Claims 1–4 in the proof of [9, Theorem 10.1], $S=\{a\in\hat{A}\mid a\sim r\ \text{for some}\ r\in R\}$ is a regular $\ast$-subring of $\hat{A}$ and there is a surjective homomorphism $g:S\to R$ such that $g(a)=r$ if and only if $a\sim r$.
Being an elementary extension of $A$, $\tilde{A}$ is directly finite and so is its subring $S$. Now, assume $sr=1$ in $R$. Consider a finite set $E\subseteq J_{0}$. According to ()(*)( ∗ ), there is $e\in J_{0}$ such that $ef=f$ and $er^{\ast}f=r^{\ast}f$ for all $f\in E$. Take $a=re$ and observe that $af=ref=rf$ and $a^{\ast}f=er^{\ast}f=r^{\ast}f$ for all $f\in E$. Thus, the set
$\Sigma(x)=\bigl{\{}[xe=re]\ \&.\ [x^{\ast}e=r^{\ast}e]\&[\exists y.\,yx=1]\mid e\in J_{0}\bigr{\}}$
of formulas with a free variables $x,y$ of type $A$ and $R$, respectively, is finitely realized in $(_{R}V_{F};A)$. Indeed, given a finite subset $\Psi$ of $\Sigma(x)$ there is finite $E\subseteq J_{0}$ containing all $f\in J_{0}$ which occur in $\Sigma(x)$; choose $e$ for $E$ as above, $g,t,u$ according to Lemma 4 below, and $x=t+1-g$, $y=u+1-g$.
By saturation, there are $a\in\tilde{A}$ and $b\in\tilde{R}$ with $ba=1$ and $a\sim r$, whence $a\in S$ and $g(a)=r$. Moreover, $a$ is left cancellable in $\tilde{R}$ whence in the subring $S$ and so a unit of $S$ by regularity. Hence, $r=g(a)$ is a unit of $R$ and $s=r^{-1}$ whence $rs=1$. ∎
Lemma 4.
Consider a regular ring $R$ with ideal $I$ such that each $eRe$, $e\in I$, is unit-regular. Then for any $r,s\in R$ with $sr=1$ and idempotent $e\in I$ there are an idempotent $g\in I$, $e\in gRg$, and $t,u\in gRg$ such that $ut=g$, $te=re$, and $et=er$.
Proof.
Following [3] we consider $R$ the endomorphism ring of a (right) $R$-module, namely $M_{R}=R_{R}$. Observe that $r$ is an injective endomorphism of $M_{R}$. Let $U=\operatorname{{\sf im}}e$, $W_{1}=U+r^{-1}(U)$, $W_{2}=r(W_{1})$; in a particular, these are submodules of $M_{R}$ and $r|W_{1}$ is an isomorphism of $W_{1}$ onto $W_{2}$. By (the proof of) [4, Lemma 2] there is an idempotent $g\in I$ such that $e,re,se\in S:=gRg$. Put $W=\operatorname{{\sf im}}g$ which is a submodule of $M_{R}$, and an $S$-module under the induced action of $S$, so that $S=\operatorname{{\sf End}}(W_{S})=\operatorname{{\sf End}}(W_{R})$.
By hypothesis, $S$ is unit-regular whence, in particular, directly finite. Due to regularity of $S$, for any $h\in S$ and $S$-linear map $\phi:hS\to W$ there is an extension $\bar{\phi}\in S$, namely $\bar{\phi}|(g-h)S=0$. Due to direct finiteness, any injective such $\phi$ has an inverse in $S$. Also, by regularity, the submodules $W_{1}=\operatorname{{\sf im}}e+\operatorname{{\sf im}}se$ and $W_{2}=\operatorname{{\sf im}}e+\operatorname{{\sf im}}re$ are of the form $W_{i}=\operatorname{{\sf im}}g_{i}$ with idempotents $g_{i}\in S$.
Let $X_{i}=\operatorname{{\sf im}}(g-g_{i})$ whence $W=W_{i}\oplus X_{i}$. Since $r|W_{1}:W_{1}\to W_{2}$ is an $S$-linear isomorphism, according to [3, Theorem 3] there is an $S$-linear isomorphism $\varepsilon:X_{1}\to X_{2}$. Put $\delta(v)=\varepsilon(v)+g_{2}(r(v))$ for $v\in X_{1}$. If $\delta(v)=w\in W_{2}$ then $\varepsilon(v)\in W_{2}\cap X_{2}$ whence $\varepsilon(v)=0$ and $v=0$; it follows that $\delta$ is an $S$-linear isomorphism of $X_{1}$ onto $Y\subseteq W$ where $Y\cap W_{2}=0$. Also, $g_{2}(\delta(v))=g_{2}(r(v))$ since $g_{2}(X_{2})=0$. Define $t\in S$ as $t(v+w)=r(v)+\delta(w)$ for $v\in W_{1}$ and $w\in X_{1}$. $t$ is injective whence it has inverse $u$ in $S$. ∎
An example of a simple regular *-ring which is not finite is obtained as follows: Let $V_{F}$ a vector space of countably infinite dimension, and $R=\operatorname{{\sf End}}(V_{F})/J(V_{F})$. Of course, $R$ is not directly finite. Define the involution on the direct product $R\times R^{op}$ by exchange: $(r,s)^{*}=(s,r)$ to obtain the *-ring $S$. Now, if $rs=1$ but $sr\neq 1$ then $xx^{*}=1$ but $x^{*}x\neq 1$ in $S$ for $x=(r,s)$.
Problem 1.
Is every simple directly finite *-regular ring representable?